[Brahms-l] Shift report 20050616 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Jun 16 2005 - 00:04:34 EDT
Shifter leader was Alex Givskov
Shifters were 


1. Machine status normal during shift.<br>

2. Run #15511 & 15512 pedestal runs.<br>
Run #15509 with FFS at 15 deg.<br>
Run #15517-15521 with FFS back at 4 deg.<br>
Run #15510,15513-15516 may be discarded due to partition fb3 was not on.<br>

3. D1 and D2 should be working properly now.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jun 16 00:04:51 2005

This archive was generated by hypermail 2.1.8 : Thu Jun 16 2005 - 00:04:58 EDT